OpenStudy (anonymous):

Gas pressure is caused by: A.) too many gas molecules in a gas tank B.) too few gas molecules in a gas tank. C.) collsions of the gas particles with the walls of the tank D.) an increase in temperature. What is the molarity of a solution prepared by dissolving 20grams of sodium hydroxide (MW = 40g/mol) in water and diluting to 1.0L? A.) 0.25M B.) 0.50M C.) 1.0M D.) 2.0M The solute in a solution... A.) is the substance being dissolved. B.) is the substance being dissolved. C.) is the solute plus the solvent. D.) is the solvent plus the solution.

OpenStudy (diogop110):

C , A, (M=n/V --> if in 40g iI have 1 mol, in 20g I have 0.5 mol. so M=0.5/1L= 0.25M, A and B, is the substance being dissolved.

OpenStudy (anonymous):

Thanks so much @diogop110

OpenStudy (diogop110):

=)

OpenStudy (anonymous):

I have another question can you help me?